OpenStudy (anonymous):

Find the distance between the points (-4, -5) and (3, -1)

OpenStudy (anonymous):

\[\sqrt{65}\] \[\sqrt{37}\] \[\sqrt{17}\]

OpenStudy (anonymous):

answers^

OpenStudy (saifoo.khan):

Do you know the distance formula?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

This will help you, http://0.tqn.com/d/math/1/0/A/1/distance.gif

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this looks easy

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

i get the the \[\sqrt{17}\]

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

RONG!

OpenStudy (anonymous):

what i followed ur formula

OpenStudy (saifoo.khan):

Check your working again..

OpenStudy (anonymous):

i type it in to my calculator with the # in the question

OpenStudy (saifoo.khan):

try again.

OpenStudy (anonymous):

just a simple fact that (-4-3)^2=49 reveals the answer.

OpenStudy (anonymous):

So now you know that the answer is \(\sqrt{65} \)

OpenStudy (anonymous):

\[a^2=4^2+7^2\] can you see this

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[a =\sqrt{4^2+7^2} =\sqrt{65}\]

OpenStudy (anonymous):

The distance formula is d=sqrt%28%28x%5B1%5D-x%5B2%5D%29%5E2%2B%28y%5B1%5D-y%5B2%5D%29%5E2%29

OpenStudy (saifoo.khan):

Nice coding^

OpenStudy (anonymous):

@nitz you lost me mate :D

OpenStudy (anonymous):

no i written the distance formula fr finding the distance between two points it changed into this form i dont know how